OpenStudy (honeymonster):

find the value of the unknown in the formula..help would be appreciated..formula posted in replies

OpenStudy (honeymonster):

\[\frac{ 4 }{ 3 }m(n-1)=p\] m=0.5 p=1.5 Find n

OpenStudy (anonymous):

(n-1)=p(3/4)/m n=p(3/4)/m+1 plug in the values